# VideoRFT: Incentivizing Video Reasoning Capability in MLLMs via Reinforced Fine-Tuning

### v3

@@ 4.1 Experimental Setup
Benchmark and Metric. Following previous works VideoR1 ; LLaVAOneVision ; TinyLLaVA-Video-R1 , we evaluate our approach on six video reasoning and understanding benchmarks: VSI-Bench VSIBench , VideoMMMU VideoMMMU , MMVU MMVU , MVBench MVBench , TempCompass TempCompass and VideoMME VideoMME , covering spatial reasoning, knowledge-intensive video QA, temporal logic, and general video understanding. Following conventions, we only use the subset of multiple-choice samples in MMVU, and VideoMME is evaluated without subtitles. Average accuracy is adopted as the evaluation metric.
@@
Implementation Details. We use Qwen2.5-VL-7B Qwen25VL as the base model and train VideoRFT on 8 NVIDIA A800 GPUs, with 80GB each. For efficiency, the video input is limited to 16 frames of resolution $128\times 28\times 28$ during training, where $28\!\times\!28$ is the size of each image patch, and $128$ denotes the number of patches. During inference, we increase the number of frames to $32$ and the resolution to $256\times 28\times 28$. For efficiency, we use a lightweight version of SigCLIP with 400M parameters in computing the semantic-consistency reward. The entire model is trained for one epoch of SFT followed by 1K steps of RL.
@@
LongVA-7B LongVA arXiv 24.06 29.2 23.9 - - 56.9 52.6
@@
LLaVA-OneVision-7B LLaVAOneVision arXiv 24.08 32.4 33.8 49.2 56.7 - 58.2
@@ 4.2 Main Result
As in Table 1, we compare VideoRFT against a variety of baselines, including proprietary models (i.e., GPT-4o gpt4o ), Open-Source MLLMs (e.g., Qwen2.5-VL Qwen25VL , VILA VILA , LongVA LongVA ), and contemporaneous models (e.g., Video-R1 VideoR1 , TinyLLaVA-Video-R1 TinyLLaVA-Video-R1 , VideoChat-R1 VideoChatR1 ).
@@ 4.3 Diagnostic Experiment
Reward Modeling. We further examine the effect of the reward defined in Eq. 6. Specifically, we compare two ablated variants: the first uses only the Format Reward and Accuracy Reward ($R\!=\!R_{f}\!+\!R_{a}$), while the second incorporates all three rewards directly ($R\!=\!R_{f}\!+\!R_{a}\!+\!R_{s}$). As seen, adding the semantic-consistency reward $R_{s}$ consistently improves performance over the first variant, validating its effectiveness. Finally, our full reward formulation, which conditionally activates $R_{s}$ via gating, achieves the best overall results. Notably, it brings substantial gains especially for video reasoning benchmarks, i.e., +2.2% on VSI-Bench, +0.9% on VideoMMMU and +3.3% on MMVU.
Training Curve. Fig. 6 shows the training dynamics of VideoRFT, including accuracy reward, response length, and semantic-consistency reward. The accuracy reward in Fig. 6 (a) shows an overall increasing trend, which means that GRPO contributes to gradually enhancing the model’s accuracy. Moreover, we observe steady improvements in response length and semantic reward. In the early stages, both metrics increase significantly, indicating that the model begins generating more complex and semantically grounded reasoning chains.
@@
(a) Accuracy reward.
(b) Response length.
(c) Semantic-consistency reward.
Figure 6: Training curves of VideoRFT in the RL phase.
@@ 5.1 Multimodal Reasoning in MLLMs
Enabling reasoning in MLLMs has become a central objective in recent research li2024enhancing ; AoTD ; MMCoT ; DDCoT . In the image domain, early works such as MMCoT MMCoT and DDCoT DDCoT disentangle perception and reasoning by treating visual understanding as input prompts for subsequent inference. In the video domain, VoT VoT and STEP STEP decompose video reasoning into predefined stages, employing template-based prompting to facilitate multi-step inference. DoraemonGPT DoraemonGPT models video understanding through symbolic memory and external tool sequences, yet still follows a modular reasoning paradigm. While these methods offer structured supervision, their rigid designs often limit generalization across diverse temporal and causal scenarios. Recently, rule-based RL has emerged as a promising paradigm for promoting multimodal reasoning in MLLMs. Pioneering efforts such as Visual-RFT VisualRFT , R1-VL R1VL , and Reason-RFT ReasonRFT directly adapt rule-based RL to image perception tasks. Follow-ups like Vision-R1 VisionR1 and R1-OneVision R1OneVision further demonstrate its effectiveness in enabling CoT reasoning on images. Concurrently, this paradigm has also been explored for video understanding VideoChatR1 ; VideoR1 ; TinyLLaVA-Video-R1 . Despite encouraging progress, these methods face a fundamental bottleneck: the lack of large-scale, high-quality video CoT datasets, which limits the full potential of RFT in the video domain. Our work addresses this gap by proposing a scalable and cognitively inspired pipeline to automatically mine high-quality CoT annotations for videos. Beyond this, we introduce a novel reward modeling strategy based on cross-modal semantic consistency, which explicitly guides MLLMs to generate visually grounded reasoning traces, and proves to be highly effective in improving model performance.
@@ 5.2 Multimodal CoT Dataset Construction
CoT has proven effective for enhancing the reasoning capabilities of LLMs by encouraging step-by-step reasoning CoT ; AutomaticCoT . Constructing high-quality CoT data in multimodal settings, particularly for video reasoning, remains a major challenge due to the temporal complexity and visual ambiguity of video data VideoEspresso ; VideoCoT ; VisualCoT . Recent works have explored CoT construction in both image and video domains. LLaVA-CoT llavacot , Vision-R1 VisionR1 , and R1-OneVision R1OneVision simply convert visual inputs to textual descriptions before reasoning. This often leads to hallucinations and weak semantic alignment. Video-R1 VideoR1 adopts a simplistic prompting strategy that encourages MLLMs to generate CoT by inserting “let me think”, “wait”, etc.into responses. However, such CoTs merely mimic the surface form of human thinking without engaging in genuine reasoning. VideoEspresso VideoEspresso generates CoT data by prompting GPT-4o with a small set of selected key frames. Due to the sparse visual context and reliance on a text-only model, the generated CoTs often lack grounding in the actual video content and are prone to hallucinations. In contrast, our CoT data combines the reasoning abilities of reasoning LLMs and the multimodal abilities of MLLMs, ensuring the reasoning depth and visual grounding of CoT data. Moreover, we use cognition-inspired prompts to enable the reasoning model to generate CoT data that is more in line with human cognition.
@@ 6 Conclusion
In this work, we introduce VideoRFT, a novel approach for incentivizing cognitive video reasoning capabilities in MLLMs through reinforced fine-tuning. To accomplish this, we propose a cross-modal pipeline that generates high-quality cognitive video CoT data simulating human reasoning processes, resulting in two large-scale, high-quality datasets: VideoRFT-CoT-102K and VideoRFT-RL-310K. Furthermore, to strengthen the RL phase, we develop semantic-consistency guided reward modeling to explicitly encourage the alignment between reasoning traces and visual evidence. Extensive experiments across six video reasoning benchmarks demonstrate that VideoRFT consistently surpasses a variety of advanced MLLMs. We expect this work to lay a foundation for future efforts in RFT-based video reasoning.

### v4

@@ 4.1 Experimental Setup
Benchmark and Metric. Following previous works VideoR1 ; LLaVAOneVision ; TinyLLaVA-Video-R1 , we evaluate our approach on six video reasoning and understanding benchmarks: VSI-Bench VSIBench , VideoMMMU VideoMMMU , MMVU MMVU , MVBench MVBench , TempCompass TempCompass , and VideoMME VideoMME , covering spatial reasoning, knowledge-intensive video QA, temporal logic, and general video understanding. Following conventions, we only use the subset of multiple-choice samples in MMVU, and VideoMME is evaluated without subtitles. Average accuracy is adopted as the evaluation metric.
@@
Implementation Details. We use Qwen2.5-VL-7B Qwen25VL as the base model and train VideoRFT on 8 NVIDIA A800 GPUs, with 80GB each. For efficiency, the video input is limited to 16 frames, with each frame processed into $128\times 28\times 28$ resolution during training, where $28\times 28$ is the size of each image patch, and $128$ denotes the number of patches. During inference, we increase the number of frames to $32$ and the resolution to $256\times 28\times 28$. For efficiency, we use a lightweight version of SigLIP with 400M parameters in computing the semantic-consistency reward. The entire model is trained for one epoch of SFT followed by 1K steps of RL.
@@
LongVA-7B LongVA TMLR 24 29.2 23.9 - - 56.9 52.6
@@
LLaVA-OneVision-7B LLaVAOneVision TMLR 24 32.4 33.8 49.2 56.7 - 58.2
@@ 4.2 Main Result
As shown in Table 1, we compare VideoRFT against a variety of baselines, including proprietary models (i.e., GPT-4o gpt4o ), Open-Source MLLMs (e.g., Qwen2.5-VL Qwen25VL , VILA VILA , LongVA LongVA ), and contemporaneous models (e.g., Video-R1 VideoR1 , TinyLLaVA-Video-R1 TinyLLaVA-Video-R1 , VideoChat-R1 VideoChatR1 ).
@@ 4.3 Diagnostic Experiment
Training Data. To assess the impact of the cross-modal refinement in Eq. 2 on data quality, we conduct SFT+RL using the data generated by Eq. 1, denoted as $\text{CoT}_{v}^{(0)}$, resulting in a variant, i.e., without CoT Refinement. As shown in Table 2, this consistently leads to performance drops across all six benchmarks, i.e., -2.3% on VSI-Bench, -3.7% on MMVU, and -7.0% on VideoMME. This demonstrates that our cross-modal refinement effectively mitigates errors and hallucinations in the initial CoTs, thereby enhancing data quality and ultimately improving model performance.
@@
Reward Modeling. We further examine the effect of the reward defined in Eq. 6. Specifically, we compare two ablated variants: the first uses only the Format Reward and Accuracy Reward ($R=R_{f}+R_{a}$), while the second incorporates all three rewards directly ($R=R_{f}+R_{a}+R_{s}$). As seen from Table 2, adding the semantic-consistency reward $R_{s}$ consistently improves performance over the first variant, validating its effectiveness. Finally, our full reward formulation, which conditionally activates $R_{s}$ via a gating mechanism (i.e., the indicator function $\mathbb{1}[R_{a}>0]$), achieves the best overall results. Notably, it brings substantial gains especially for video reasoning benchmarks, i.e., +2.2% on VSI-Bench, +3.3% on MMVU, and +3.5% on VideoMME.
@@ 5.1 Multimodal Reasoning in MLLMs
Enabling reasoning in MLLMs has become a central objective in recent research li2024enhancing ; AoTD ; MMCoT ; DDCoT . In the image domain, early works such as MMCoT MMCoT and DDCoT DDCoT disentangle perception and reasoning by treating visual understanding as input prompts for subsequent inference. In the video domain, VoT VoT and STEP STEP decompose video reasoning into predefined stages, employing template-based prompting to facilitate multi-step inference. DoraemonGPT DoraemonGPT models video understanding through symbolic memory and external tool sequences, yet still follows a modular reasoning paradigm. While these methods offer structured supervision, their rigid designs often limit generalization across diverse temporal and causal scenarios. Recently, rule-based RL has emerged as a promising paradigm for promoting multimodal reasoning in MLLMs. Pioneering efforts such as Visual-RFT VisualRFT , R1-VL R1VL , and Reason-RFT ReasonRFT directly adapt rule-based RL to image perception tasks. Follow-ups like Vision-R1 VisionR1 and R1-OneVision R1OneVision further demonstrate its effectiveness in enabling CoT reasoning on images. Concurrently, this paradigm has also been explored for video understanding VideoR1 ; VideoChatR1 ; TinyLLaVA-Video-R1 . Despite encouraging progress, these methods face a fundamental bottleneck: the lack of large-scale, high-quality video CoT datasets, which limits the full potential of RFT in the video domain. Our work addresses this gap by proposing a scalable and cognitively inspired pipeline to automatically mine high-quality CoT annotations for videos. Beyond this, we introduce a novel reward modeling strategy based on cross-modal semantic consistency, which explicitly guides MLLMs to generate visually grounded reasoning traces, and proves to be highly effective in improving model performance.
@@ 5.2 Multimodal CoT Dataset Construction
CoT has proven effective for enhancing the reasoning capabilities of LLMs by encouraging step-by-step reasoning CoT ; AutomaticCoT . Constructing high-quality CoT data in multimodal settings, particularly for video reasoning, remains a major challenge due to the temporal complexity and visual ambiguity of video data VideoEspresso ; VisualCoT ; VideoCoT . Recent works have explored CoT construction in both image and video domains. LLaVA-CoT llavacot , Vision-R1 VisionR1 , and R1-OneVision R1OneVision simply convert visual inputs to textual descriptions before reasoning. This often leads to hallucinations and weak semantic alignment. Video-R1 VideoR1 adopts a simplistic prompting strategy that encourages MLLMs to generate CoT by inserting “let me think”, “wait”, etc.into responses. However, such CoTs merely mimic the surface form of human thinking without engaging in genuine reasoning. VideoEspresso VideoEspresso generates CoT data by prompting GPT-4o with a small set of selected key frames. Due to the sparse visual context and reliance on a text-only model, the generated CoTs often lack grounding in the actual video content and are prone to hallucinations. In contrast, our CoT data combines the reasoning abilities of reasoning LLMs and the multimodal abilities of MLLMs, ensuring the reasoning depth and visual grounding of CoT data. Moreover, we use cognition-inspired prompts to enable the reasoning model to generate CoT data that is more in line with human cognition.
@@ 6 Conclusion
In this work, we introduce VideoRFT, a novel approach for incentivizing cognitive video reasoning capabilities in MLLMs through reinforced fine-tuning. To accomplish this, we propose a cross-modal pipeline that generates high-quality cognitive video CoT data simulating human reasoning processes, resulting in two large-scale datasets: VideoRFT-CoT-102K and VideoRFT-RL-310K. Furthermore, to strengthen the RL phase, we develop semantic-consistency guided reward to explicitly encourage the alignment between reasoning traces and visual evidence. Extensive experiments across six benchmarks demonstrate that VideoRFT consistently surpasses a variety of advanced MLLMs. We expect this work to lay a foundation for future efforts in RFT-based video reasoning.